OpenStudy (yacoub1993):

Find the quotient z1/z2 of the complex numbers. Leave answer in polar form. z1= sqrt3 (cos 7pie/4 + i sin 7pie/4) z2 = sqrt6 (cos 9pie/4 + i sin 9pie/4)

OpenStudy (yacoub1993):

@amistre64 Find the quotient z1/z2 of the complex numbers. Leave answer in polar form. z1= sqrt3 (cos 7pie/4 + i sin 7pie/4) z2 = sqrt6 (cos 9pie/4 + i sin 9pie/4)

OpenStudy (yacoub1993):

@phi Find the quotient z1/z2 of the complex numbers. Leave answer in polar form. z1= sqrt3 (cos 7pie/4 + i sin 7pie/4) z2 = sqrt6 (cos 9pie/4 + i sin 9pie/4)

OpenStudy (phi):

Did they teach you the method to do this ?

OpenStudy (yacoub1993):

no actually i am taching miself through your help openstudy Could you help me out

OpenStudy (phi):

because I would write both in the form \( A e^{i \theta} \)

OpenStudy (yacoub1993):

can you show me how to do it

OpenStudy (phi):

Euler's formula http://en.wikipedia.org/wiki/Euler%27s_formula \[ e^{ ix} = \cos x + i \sin(x) \] so \[ z_1= \sqrt{3} \left(\cos\left( \frac{7\pi}{4}\right) + i \sin\left( \frac{7\pi}{4}\right)\right) = \sqrt{3} e^{i \frac{7\pi}{4}}\] and \[ z_2= \sqrt{6} e^{i \frac{9\pi}{4}}\] so \[ \frac{z_1}{z_2}= \frac{\sqrt{3} e^{i \frac{7\pi}{4}}}{\sqrt{6} e^{i \frac{9\pi}{4}}} \]

OpenStudy (yacoub1993):

i have a question @phi What does "i" stand for?

OpenStudy (phi):

i = \( \sqrt{-1} \) as you know, there is no "normal" i.e. real number that when you square it, gives -1. 1*1= 1 and -1 * -1 = 1 so people said, let's just say there is such a number (and because sqrt(-1) is awkward to write out all the time) they said, we will call it i i*i= -1 by definition You might be able to follow Khan's great video on this https://www.khanacademy.org/math/calculus/sequences_series_approx_calc/maclaurin_taylor/v/euler-s-formula-and-euler-s-identity

OpenStudy (phi):

to finish, use the "exponent rule" \[ \frac{a^b}{a^c} = a^{b-c} \] \[ \frac{z_1}{z_2}= \frac{\sqrt{3} e^{i \frac{7\pi}{4}}}{\sqrt{6} e^{i \frac{9\pi}{4}}} = \frac{\sqrt{3}}{\sqrt{6}} e^{- \frac{\pi}{2}}= \frac{\sqrt{2}}{2}\left( \cos\left(- \frac{\pi}{2}\right) +i \sin\left(- \frac{\pi}{2}\right)\right) \] we can simplify that using cos(-x)= cos(x) and sin(-x)= - sin(x) \[ \frac{\sqrt{2}}{2}\left( \cos\left(\frac{\pi}{2}\right) -i \sin\left(\frac{\pi}{2}\right)\right) \] of course we know cos(pi/2) is 0 and sin(pi/2) = 1 so we can simplify further